# M. Belianovich, G. E. Fasshauer, A. Narayan, and V. Shankar

### v2

@@ 3.2.1 $f(x,y)=(x^{2}+y^{2})^{\frac{3}{2}}$
Figure 4 (left) shows that the relative $\ell_{2}$ error decays faster for the unified interpolant with $\epsilon=10$ than for either the unified interpolant in the polynomial limit (Diag) or standard polynomial least squares (PLS).
@@
We now test the hybrid interpolant on the target function $f(x,y)=\exp\left(\frac{(x+y)^{2}}{0.2}\right)$, once again on the unit disk $D$. Since this target is analytic, we expect rapid convergence rates for all our methods. Figure 4 (right) shows that all three methods reach a minimum error of $O(10^{-15})$ at the same rates. Interestingly, it appears that Wendland kernels do not significantly affect the approximation errors when the target function is sufficiently smooth; this is apparent when contrasting with the convergence results for $f(x)=(x^{2}+y^{2})^{\frac{3}{2}}$, where the unified interpolant in the non-diagonal limit is clearly superior.
@@ 3.3.1 $f(x,y,z)=(x^{2}+y^{2}+z^{2})^{\frac{3}{2}}$
Figure 5 (left) shows that the relative $\ell_{2}$ error for this unified interpolant decays at roughly the same rate for $C^{2}(\mathbb{R}^{3})$ kernel as in the PLS or Diag approximations. However on the finest node set, the unified interpolant attains an error that is half an order of magnitude smaller. We see that the same trend as in the 2D case: the unified interpolant is superior to standard polynomial least squares for rough target functions on Euclidean domains.
@@
We also tested the unified interpolant on the analytic target function $f(x,y,z)=\exp\left(\frac{(x+y+z)^{2}}{0.8}\right)$ on the unit ball. In this case, one expects rapid convergence for all approximations. Figure 5 (right) verifies that the error decays rapidly for all three approximations. The unified interpolant appears to perform very slightly better on the finest node set, but the differences are nowhere as drastic as for rougher targets. The takeway appears to be that the unified interpolant is superior for rougher target functions on Euclidean domains.
@@ 3.4 Results on $\mathbb{M}\subset\mathbb{R}^{3}$
Figure 6: Convergence for a $C^{1}(\mathbb{M})$ target function on the sphere $\mathbb{S}^{2}$ (left) and torus $\mathbb{T}$ (right) as a function of $\sqrt{N}$. The top row shows the target function $f(x,y,z)=x^{2}|x|+y^{2}|y|+z^{2}|z|$ visualized on each manifold and the bottom row shows the convergence plots.
We also present results on manifolds. Our goal is to demonstrate feasibility, rather than an extensive study. While the kernel Gramian $A$ is itself full-rank on a algebraic manifold for distinct data sites $X$, the polynomial matrix $P$ is rank-deficient when the zero-locus of the polynomial coincides with the algebraic variety describing the manifold itself. As a consequence, the matrix $B=L^{-1}P$ is rank-deficient also.
@@
However, as mentioned previously, this can be rectified with a column-pivoted QR factorization of $B$, which also returns a pivot indexing matrix $E$. We select a truncation tolerance as $\tau=\max([N,M]{\rm eps}(\|R\|_{\infty})$. Then, we set $r$ to be the last entry of the diagonal of $R$ that is greater than the tolerance (assuming the diagonal entries are sorted in descending order). We set $\tilde{Q}=Q(:,1:r)$, the first $r$ columns of $Q$, and $\tilde{R}=R(1:r,1:r)$. This lets us safely solve for $r$ polynomial coefficients $\tilde{\mathbf{d}}$ in the usual way using $\tilde{\mathbf{d}}=\tilde{R}^{-1}\tilde{Q}^{T}\mathbf{g}$. Finally, since we require $M$ coefficients, we set $\mathbf{d}(E)=\tilde{\mathbf{d}}$ with the remaining entries being set to zero. This minimally intrusive approach to rectifying the rank-deficiency of $B$ allows us to continue to use our efficient numerical linear algebra. $\bm{c}$ is then computed as in the Euclidean case.
@@
Having observed no differences between the different approaches for smooth target functions on manifolds, we test our approach on the $C^{1}(\mathbb{M})$ target function $f(x,y,z)=x^{2}|x|+y^{2}|y|+z^{2}|z|$ on the unit sphere $\mathbb{S}^{2}$ and the torus $\mathbb{T}=\bigl{\{}(x,y,z)\in\mathbb{R}^{3}:\left(\sqrt{x^{2}+y^{2}}-R^{2}\right)^{2}+z^{2}=r^{2}\bigr{\}}$ with radii $R=1$ and $r=1/3$, restricting ourselves to the $C^{2}(\mathbb{R}^{3})$ Wendland function and Legendre polynomials chosen in $\mathbb{R}^{3}$. On the sphere, we used generalized spiral points [25, 26, 35] and on the torus, we use the staggered hexagonal nodes first used in [34], which also have similar properties. In both cases, we set the shape parameter to $\epsilon=7$ corresponding to a condition number of $K_{t}=O(10^{3})$ on the finest node set, which results in increasing density in the kernel Gramian $A$ as $N$ is increased. The results are shown in Figure 6 (bottom row). Interestingly, despite the function being rough, the unified interpolant underperforms the PLS approximation slightly, though both appear to converge at the same rate. The intuition from the Euclidean case fails.
@@
=$\mathbb{H}=\bigl{\{}(x,y,z)\in\mathbb{R}^{3}:x^{2}+y^{2}+z^{2}=1,\;z\geq 0\bigr{\}},$
@@
We generate $N$ “almost” Fibonacci-spiral nodes with mild clustering toward the equator $z=0$. Given a clustering exponent $q>1$ and for $k=0,\dots,N-1$ we set
@@
This places $(x_{k},y_{k},z_{k})$ quasi-uniformly in area over $\mathbb{M}=\mathbb{H}$, with increased density near the boundary circle $z=0$; these nodes are shown in Figure 7 (left). We then conducted a convergence study on the same target function used on $\mathbb{S}^{2}$ and $\mathbb{T}$. The results along are shown in Figure 7 (right).
Figure 7: Convergence for a $C^{1}(\mathbb{M})$ target function on the upper hemisphere as a function of $\sqrt{N}$. We show a sample node set (left) and the actual convergence results (right).
@@
Note: Upon seeing these results, we returned to approximation on $[-1,1]$ and attempted to use data sites that were less clustered than Chebyshev extrema to see if the Wendland kernels had a similar effect there. We accomplished this by using the Kosloff-Tal-Ezer map $x_{k}=\frac{\arcsin(\alpha x_{k}^{\rm cheb})}{\arcsin(\alpha)}$ [18] for various values of $\alpha$, but we saw no differences from the results in Section 3.1 beyond those caused by instabilities from insufficient clustering. It appears unified interpolation is mainly useful in $d=2$ and higher.
@@ 4 Summary and Future Work
Interestingly, similar observations have been made about PHS kernels in conjunction with polynomials, albeit in the context of local interpolation. We believe our unified interpolation scheme fits into the body of work on PHS kernels with polynomials, but also into other work such as universal kriging [20]; some simple interpolation examples using Wendland kernels with additional polynomials, i.e., univeral kriging, can be found in [7, Sect. 17.3].

### v3

@@ 3.2.1 $f(x,y)=(x^{2}+y^{2})^{\frac{3}{2}}$
Figure 6 (left) shows that the relative $\ell_{2}$ error decays faster for the unified interpolant with $\epsilon=10$ than for either the unified interpolant in the polynomial limit (Diag) or standard polynomial least squares (PLS), while PHS+poly is uniformly less accurate and shows mild deterioration at the finest grid.
@@
We now test the hybrid interpolant on the target function $f(x,y)=\exp\left(\frac{(x+y)^{2}}{0.2}\right)$, once again on the unit disk $D$. Since this target is analytic, we expect rapid convergence rates for all our methods.
Figure 6 (right) shows that three methods reach a minimum error of $O(10^{-15})$ at the same rates, whereas PHS+poly remains several orders of magnitude less accurate over the whole range. Interestingly, it appears that Wendland kernels do not significantly affect the approximation errors when the target function is sufficiently smooth; this is apparent when contrasting with the convergence results for $f(x)=(x^{2}+y^{2})^{\frac{3}{2}}$, where the unified interpolant in the non-diagonal limit is clearly superior.
Figure 7: Computational costs vs. $N^{1/2}$ for $f(x,y)=(x^{2}+y^{2})^{\frac{3}{2}}$ (top row) and $f(x,y)=\exp\left(\frac{(x+y)^{2}}{0.2}\right)$ (bottom row) using the $C^{2}(\mathbb{R}^{3})$ Wendland function. Left: assembly and solve time; right: evaluation time. Curves shown for PLS, Diag, and FS with fixed shape parameter $\epsilon=10$.
3.2.3 Timings on the disk.
We next report computational cost as a function of $N^{1/2}$ on the same boundary–clustered disk node sets. Figure 7 presents two panels per target: assembly and solve time on the left and evaluation time on the right, for the three approximations (PLS, Diag, and FS with fixed $K_{t}=O(10^{3})$). Two robust, dataset–independent trends are evident. First, the evaluation curves for PLS, Diag, and FS are nearly indistinguishable across all grid sizes; they increase smoothly with $N^{1/2}$ and track each other closely. In contrast, PHS+poly forms an upper bundle in the right column, evaluating consistently more slowly due to dense global kernel evaluation. Second, the separation appears in assembly and solve: as $N^{1/2}$ grows, the three methods no longer behave the same, with one method clearly more expensive.
The ordering in assembly and solve is consistently $\text{PLS}\approx\text{Diag}\ll\text{FS}$. This gap stems from how FS achieves the target condition number: the support sizes selected to enforce $K_{t}$ produce denser local stencils and a larger sparse system, and they incur additional neighbor–search and Wendland Gramian block–assembly overhead. As the grids refine, these effects accumulate and the FS curve acquires a noticeably steeper slope than PLS or Diag. Augmenting with PHS+poly, the ordering becomes $\text{PLS}\approx\text{Diag}\ll\text{FS}\ll\text{PHS+poly}$ in the top panels: the dense saddle point solve makes PHS+poly the most expensive and gives it the steepest growth with $N^{1/2}$. This pattern holds for both targets, $f(x,y)=(x^{2}+y^{2})^{\frac{3}{2}}$ and $\exp\!\big(\frac{(x+y)^{2}}{0.2}\big)$, indicating that the premium is driven by support selection (FS) and global density (PHS+poly), rather than by target smoothness.
@@ 3.3.1 $f(x,y,z)=(x^{2}+y^{2}+z^{2})^{\frac{3}{2}}$
Figure 8 (left) shows that the relative $\ell_{2}$ error for this unified interpolant decays at roughly the same rate for $C^{2}(\mathbb{R}^{3})$ kernel as in the PLS or Diag approximations. Interestingly, PHS+poly is the most accurate across the tested sizes and improves monotonically. In contrast, unified/PLS/Diag uniformly show higher errors. However on the finest node set, the unified interpolant attains an error that is half an order of magnitude smaller than the PLS or Diag methods. We see that the same trend as in the 2D case: the unified interpolant is superior to standard polynomial least squares for rough target functions on Euclidean domains.
@@
We also tested the unified interpolant on the analytic target function $f(x,y,z)=\exp\left(\frac{(x+y+z)^{2}}{0.8}\right)$ on the unit ball. In this case, one expects rapid convergence for all approximations. Figure 8 (right) verifies that the error decays rapidly for other three approximations, while PHS+poly remains orders of magnitude less accurate over the entire range. The unified interpolant appears to perform very slightly better on the finest node set, but the differences are nowhere as drastic as for rougher targets. The takeaway appears to be that the unified interpolant is superior for rougher target functions on Euclidean domains.
Figure 9: Computational costs vs. $N^{1/3}$ for $f(x,y,z)=(x^{2}+y^{2}+z^{2})^{\frac{3}{2}}$ (top row) and $f(x,y,z)=\exp\left(\frac{(x+y+z)^{2}}{0.8}\right)$ (bottom row) using the $C^{2}(\mathbb{R}^{3})$ Wendland function. Left: assembly and solve time; right: evaluation time. Curves shown for PLS, Diag, and FS with fixed shape parameter $\epsilon=5$.
3.3.3 Timings on the unit ball.
We next measure computational cost versus $N^{1/3}$ on boundary-clustered node sets in the unit ball. Figure 9 again shows assembly and solve (left column) and evaluation time (right column) for PLS, Diag, and FS at fixed $K_{t}=O(10^{3})$. The global picture mirrors the unit disk case (2D). Evaluation curves for all three methods are tightly clustered and grow smoothly with $N^{1/3}$, reflecting that once coefficients are computed, the same hybrid basis is evaluated. In contrast, the assembly and solve curves separate as $N^{1/3}$ increases. Including PHS+poly, the right column exhibits two evaluation tiers: PLS/Diag/FS occupy the faster tier with smooth growth, whereas PHS+poly is consistently slower due to dense global kernel evaluation (Figure 9).
As in 2D, the assembly and solve ordering is $\text{PLS}\approx\text{Diag}\ll\text{FS}$. The additional cost for FS is explained by the larger supports required to enforce the target condition number, which increase stencil density and system size, together with the overhead of neighbor searches and Wendland Gramian construction. These factors yield a steeper FS slope and a widening gap on the finer 3D grids. Augmenting with PHS+poly sharpens the contrast: in the top panels the ordering becomes $\text{PLS}\approx\text{Diag}\ll\text{FS}\ll\text{PHS+poly}$, with PHS+poly the most expensive and exhibiting the steepest growth with $N^{1/3}$ (Figure 9). The effect is visible for both $f(x,y,z)=(x^{2}+y^{2}+z^{2})^{3/2}$ and the analytic $\exp\!\big(\frac{(x+y+z)^{2}}{0.8}\big)$, confirming that the increased cost is driven by support selection (FS) and global density (PHS+poly), rather than by target smoothness.
Takeaways.
Taken together, these comparisons point to a clear practical conclusion. The unified interpolant is the most effective default across our Euclidean tests: in 1D and 2D it matches, and often slightly improves upon, PLS/Diag in accuracy while retaining sparse assembly, and in 3D it remains competitive in accuracy at a fraction of the cost of global PHS+poly. The sole exception we observed is the rough, radially symmetric target $(x^{2}+y^{2}+z^{2})^{3/2}$ in 3D, for which PHS+poly delivers the smallest error. Even there, however, the advantage comes with substantially higher assembly time, slower evaluation, and a heavier memory footprint that precludes the two finest grids. For analytic targets in 3D, unified/PLS/Diag are both more accurate and markedly cheaper than PHS+poly. Throughout, evaluation costs for unified/PLS/Diag form a consistently lower band, while assembly costs rank as PLS $\approx$ Diag << FS $\ll$ PHS+poly. For these reasons, we advocate for the unified interpolant over global PHS+poly for the regimes considered here.
@@ 3.4 Results on $\mathbb{M}\subset\mathbb{R}^{3}$
Figure 10: Convergence for a $C^{1}(\mathbb{M})$ target function on the sphere $\mathbb{S}^{2}$ (left) and torus $\mathbb{T}$ (right) as a function of $\sqrt{N}$. The top row shows the target function $f(x,y,z)=x^{2}|x|+y^{2}|y|+z^{2}|z|$ visualized on each manifold. The bottom row reports relative $\ell_{2}$ error for PLS, FS, and PHS+poly (with $m=1$). On both manifolds FS and PLS are nearly coincident, while PHS+poly ($m=1$) is markedly more accurate across all $N$.
We also present results on manifolds. Our goal is to demonstrate feasibility, rather than an extensive study. Alongside FS and PLS, we include the globally supported polyharmonic spline with polynomial augmentation (PHS+poly). As mentioned previously, we take $\phi_{p}(r)=r^{3}$ and solve the standard saddle point system with a small pseudoinverse tolerance ($10^{-11}$) using the same ambient total-degree polynomial space $\mathcal{P}_{\ell}(\mathbb{R}^{3})$.
@@
The unified interpolant allows for a linear solve with greater efficiency than PHS+poly. While the Wendland kernel Gramian $A$ is itself full-rank on an algebraic manifold for distinct data sites $X$, the polynomial matrix $P$ is rank-deficient when the zero-locus of the polynomial coincides with the algebraic variety describing the manifold itself. As a consequence, the matrix $B=L^{-1}P$ is rank-deficient also. However, as mentioned previously, this can be rectified with a column-pivoted QR factorization of $B$, which also returns a pivot indexing matrix $E$. We select a truncation tolerance as $\tau=\max([N,M]{\rm eps}(\|R\|_{\infty})$. Then, we set $r$ to be the last entry of the diagonal of $R$ that is greater than the tolerance (assuming the diagonal entries are sorted in descending order). We set $\tilde{Q}=Q(:,1:r)$, the first $r$ columns of $Q$, and $\tilde{R}=R(1:r,1:r)$. This lets us safely solve for $r$ polynomial coefficients $\tilde{\mathbf{d}}$ in the usual way using $\tilde{\mathbf{d}}=\tilde{R}^{-1}\tilde{Q}^{T}\mathbf{g}$. Finally, since we require $M$ coefficients, we set $\mathbf{d}(E)=\tilde{\mathbf{d}}$ with the remaining entries being set to zero. This minimally intrusive approach to rectifying the rank-deficiency of $B$ allows us to continue to use our efficient numerical linear algebra. $\bm{c}$ is then computed as in the Euclidean case.
Having observed no differences between the different approaches for smooth target functions on manifolds, we test our approach on the $C^{1}(\mathbb{M})$ target function $f(x,y,z)=x^{2}|x|+y^{2}|y|+z^{2}|z|$ on the unit sphere $\mathbb{S}^{2}$ and the torus
$\mathbb{T}=\bigl\{(x,y,z)\in\mathbb{R}^{3}:\bigl(\sqrt{x^{2}+y^{2}}-R\bigr)^{2}+z^{2}=r^{2}\bigr\},$
@@
with radii $R=1$ and $r=1/3$, restricting ourselves to the $C^{2}(\mathbb{R}^{3})$ Wendland function and Legendre polynomials chosen in $\mathbb{R}^{3}$. On the sphere, we used generalized spiral points [28, 29, 39] and on the torus, we use the staggered hexagonal nodes first used in [35], which also have similar properties. In both cases, we set the shape parameter to $\epsilon=7$ corresponding to a condition number of $K_{t}=O(10^{3})$ on the finest node set, which results in increasing density in the kernel Gramian $A$ as $N$ is increased. The results are shown in Figure 10 (bottom row). Interestingly, despite the function being rough, the unified interpolant underperforms the PLS approximation slightly, though both appear to converge at the same rate. The intuition from the Euclidean case fails.
On the sphere $\mathbb{S}^{2}$ and torus $\mathbb{T}$ with radii $R=1$ and $r=1/3$, using $C^{2}(\mathbb{R}^{3})$ Wendland CSRBFs and ambient Legendre polynomials, FS and PLS track each other closely across all $N$ (bottom row of Fig. 10). In contrast, PHS+poly with attains substantially smaller errors—often by one to several orders of magnitude. Thus, on closed manifolds the unified interpolant behaves like its polynomial limit (PLS), while PHS+poly delivers the best accuracy.
@@
$\mathbb{H}=\bigl\{(x,y,z)\in\mathbb{R}^{3}:x^{2}+y^{2}+z^{2}=1,\;z\geq 0\bigr\},$
@@
We generate $N$ “almost” Fibonacci‐spiral nodes with mild clustering toward the equator $z=0$. Given a clustering exponent $q>1$ and for $k=0,\dots,N-1$ we set
@@
This places $(x_{k},y_{k},z_{k})$ quasi‐uniformly in area over $\mathbb{M}=\mathbb{H}$, with increased density near the boundary circle $z=0$; these nodes are shown in Figure 11 (left). We then conducted a convergence study on the same target function used on $\mathbb{S}^{2}$ and $\mathbb{T}$. The results along are shown in Figure 11 (right).
Figure 11: Convergence for a $C^{1}(\mathbb{M})$ target function on the upper hemisphere as a function of $\sqrt{N}$. We show a sample node set (left) and the relative $\ell_{2}$ error for PLS, FS, and PHS+poly (with $p=3$) vs $N^{1/2}$ (right).
@@
Note: Upon seeing these results, we returned to approximation on [1,1][-1,1] and attempted to use data sites that were less clustered than Chebyshev extrema to see if the Wendland kernels had a similar effect there. We accomplished this by using the Kosloff-Tal-Ezer map $x_{k}=\frac{\arcsin(\alpha x_{k}^{\rm cheb})}{\arcsin(\alpha)}$ [21] for various values of $\alpha$, but we saw no differences from the results in Section 3.1 beyond those caused by instabilities from insufficient clustering. It appears unified interpolation is mainly useful in $d=2$ and higher.
3.5 Ablation: isolating edge effects.
Figure 12: Ablation: boundary vs. interior (FS). Relative $l_{2}$ error on the full domain (blue) and on the interior-only set (orange), together with the edge-gain ratio $E_{\text{full}}/E_{\text{int}}$ (yellow) versus $N^{1/2}$ on the disk. Curves are representative across all targets listed in the legend; the overlap illustrates that boundary effects are modest once the support is fixed at the high-stability setting.
We quantify the boundary contribution to the error by comparing the relative $\ell_{2}$ error on (i) the full evaluation set (interior ++ boundary) and (ii) an interior-only set obtained by removing a thin boundary strip of fixed physical width. We report the edge-gain ratio
$G_{\mathrm{edge}}\;=\;\frac{E_{\mathrm{full}}}{E_{\mathrm{int}}},$
so $G_{\mathrm{edge}}\!\approx\!1$ means boundary points do not materially change the error (boundary-neutral), $G_{\mathrm{edge}}\!>\!1$ indicates boundary inflation, and $G_{\mathrm{edge}}\!<\!1$ indicates the full-domain error is slightly lower than the interior-only error.
Experimental details match the 2D disk tests: Wendland $C^{2}(\mathbb{R}^{3})$ RBFs, polynomial degree $\ell=\lfloor 0.8\,N^{1/2}\rfloor$, fixed-support (FS) baseline with shape $\epsilon=10$, boundary-clustered disk node families, and evaluation on the large reference set from DiskPoissonNodesLarge (the accompanying node set). We consider five targets with distinct edge behavior:
$f_{1}(x,y)$ $=\;\exp\!\left(-\frac{1-r}{\tau_{\mathrm{BL}}}\right),$ $f_{2}(x,y)$ $=\;\bigl(1-r\bigr)_{+}^{1/4},$
$f_{3}(x,y)$ $=\;\exp(x+y),$ $f_{4}(x,y)$ $=\;(1+9r^{2})^{-1},$
$f_{5}(x,y)$ $=\;r^{3},$ $\text{with }r=\sqrt{x^{2}+y^{2}},\quad(t)_{+}=\max(t,0).$
Figure 12 shows that for all five targets the full-domain and interior-only errors track closely as $N^{1/2}$ grows, and $G_{\mathrm{edge}}$ remains near unity (often slightly below). Thus, once FS is set to the high-stability regime, its error is boundary-neutral: the boundary layer does not degrade accuracy and can even yield a small net reduction (via locally well-conditioned Wendland Gramians and boundary clustering).
The comparison above isolates the behavior of FS itself. Our broader claim that the unified interpolant pure polynomials on bounded domains because polynomials are edge-sensitive is supported by the cross-method accuracy results (PLS vs. FS) on the disk and by the periodic (no-boundary) control, where the gap collapses. In other words, FS is boundary-neutral (as shown here), whereas global polynomials exhibit boundary amplification; this contrast explains the FS–vs–PLS accuracy gap on bounded domains and its disappearance on periodic geometries.
@@ 4 Summary and Future Work
Interestingly, similar observations have been made about PHS kernels in conjunction with polynomials, albeit in the context of local interpolation. We believe our unified interpolation scheme fits into the body of work on PHS kernels with polynomials, but also into other work such as universal kriging [23]; some simple interpolation examples using Wendland kernels with additional polynomials, i.e., univeral kriging, can be found in [10, Sect. 17.3].